Solve for the electric flux that acts on a uniform electric field with a magnitude of 8000N/C approximately passes through a 10m^2 flat surface area.


1
Expert's answer
2022-02-28T10:37:58-0500

Answer

the electric flux

"\phi=EAcos\theta\\=8000*10*cos0°\\=80000Wb"